OpenStudy (monica0597):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = x + 2. y = 4x - 1 y = 4x - 3 y = x - 3 y = x - 1

OpenStudy (anonymous):

ummmmmmmmmmmmmmmm................ lol idfk i hate math

OpenStudy (monica0597):

u think i like math? i dont but i need to do this so i can make up for almost failing :/

OpenStudy (lgbasallote):

a toast for us people who hate math. hear, hear!

OpenStudy (anonymous):

i never said u did , but ur better than i am ,

OpenStudy (monica0597):

^ not really, i suck at this. i know some things but not al. exhibit A this question

OpenStudy (monica0597):

all*

OpenStudy (anonymous):

lol , im better at geometry than algebra , and thats how far it gets

OpenStudy (monica0597):

i haven't even taken geometry yet. i start taking it this year smh

OpenStudy (anonymous):

okay here are the cues you need to understand slope, slope form, and point-slope formula

OpenStudy (anonymous):

really... slope intercept form -- y=mx+b m is slope and b is the y intercept for parallel lines- slope of both lines must be equal and the y intercept must be different so since you're given y=x+2 you can easily find slope just substitute the slope and point into point intercept form-- y-y1=m(x-x1) then simplify to slope intercept form

OpenStudy (monica0597):

thanks

OpenStudy (anonymous):

would you like an e-book with the suggestions?

OpenStudy (monica0597):

sure,why not ?

OpenStudy (anonymous):

http://finedrafts.com/files/Larson%20PreCal%208th/Larson%20Precal%20CH1.pdf read beginning section 1.3, page 24

OpenStudy (monica0597):

thank u!

OpenStudy (anonymous):

don't thank me yet.

OpenStudy (anonymous):

c and d are both parrallel but none passed through that point

OpenStudy (monica0597):

thanks to all of u :)